Find how long it takes for $1500 to double if it is invested at 7% interest compounded monthly. Use the formula A=P(1+r/n)^nt
Answer by stanbon(75887)

Find how long it takes for $1500 to double if it is invested at 7% interest compounded monthly. Use the formula A=P(1+r/n)^nt
--------
3000 = 1500(1+(0.07/12))^(12*t)
------
(1.0058)^12t = 2
------
12t*log(1.0058) = log(2)
-------
12t = log(2)/log(1.0058)
-------
12t = 119.85
-----
time = 10 years when rounded up
==========
Cheers,
Stan H.
==========